# WAFT: Warping-Alone Field Transforms for Optical Flow

### v1

@@ Appendix A Appendix
Sequence WAFT (Ours) DPFlow [34] Flowformer++ [44] FlowDiffuser [26] DDVM [43]
Perturbed Market 3 0.953 (0.416) 0.892 (0.423) 0.958 (0.511) 0.897 (0.490) 0.787 (0.372)
Perturbed Shaman 1 0.208 (0.184) 0.213 (0.201) 0.251 (0.236) 0.267 (0.255) 0.219 (0.196)
Ambush 1 19.40 (3.146) 8.366 (2.970) 6.610 (2.605) 7.224 (2.583) 29.33 (14.07)
Ambush 3 3.669 (2.822) 3.019 (1.729) 2.939 (1.816) 3.148 (1.828) 2.855 (3.016)
Bamboo 3 0.466 (0.446) 0.486 (0.438) 0.546 (0.513) 0.594 (0.508) 0.415 (0.380)
Cave 3 2.260 (1.574) 2.341 (1.631) 2.344 (1.477) 2.464 (1.433) 2.042 (1.658)
Market 1 0.851 (0.434) 0.890 (0.491) 1.073 (0.550) 1.238 (0.517) 0.719 (0.467)
Market 4 6.652 (3.888) 7.939 (3.834) 8.086 (4.450) 8.024 (4.027) 5.517 (3.971)
Mountain 2 0.175 (0.092) 0.177 (0.078) 0.288 (0.118) 0.409 (0.101) 0.176 (0.095)
Temple 1 0.517 (0.332) 0.511 (0.302) 0.657 (0.359) 0.789 (0.340) 0.452 (0.284)
Tiger 0.503 (0.395) 0.571 (0.411) 0.595 (0.430) 0.636 (0.391) 0.413 (0.344)
Wall 2.038 (0.850) 2.011 (1.231) 1.727 (0.793) 1.692 (0.734) 1.639 (3.210)
All 2.337 (1.094) 1.975 (1.046) 1.943 (1.073) 2.026 (1.016) 2.475 (1.754)
Table 6: We report the endpoint-error (EPE) on all sequences of Sintel [3], shown in the format “final-epe (clean-epe)”. We highlight the best result on each sequence.
@@
We show the sequence results of several representative methods on Sintel [3] in Table 6.

### v2

@@ Appendix A Appendix
Sequence WAFT-Twins-a2 DPFlow Flowformer++ FlowDiffuser DDVM SAMFlow
Perturbed Market 3 0.893 (0.460) 0.892 (0.423) 0.958 (0.511) 0.897 (0.490) 0.787 (0.372) 0.932 (0.493)
Perturbed Shaman 1 0.174 (0.163) 0.213 (0.201) 0.251 (0.236) 0.267 (0.255) 0.219 (0.196) 0.241 (0.219)
Ambush 1 20.561 (2.526) 8.366 (2.970) 6.610 (2.605) 7.224 (2.583) 29.33 (14.07) 10.586 (2.733)
Ambush 3 3.173 (2.048) 3.019 (1.729) 2.939 (1.816) 3.148 (1.828) 2.855 (3.016) 3.411 (1.779)
Bamboo 3 0.460 (0.423) 0.486 (0.438) 0.546 (0.513) 0.594 (0.508) 0.415 (0.380) 0.522 (0.473)
Cave 3 2.199 (1.567) 2.341 (1.631) 2.344 (1.477) 2.464 (1.433) 2.042 (1.658) 2.475 (1.445)
Market 1 0.851 (0.384) 0.890 (0.491) 1.073 (0.550) 1.238 (0.517) 0.719 (0.467) 1.060 (0.491)
Market 4 7.513 (3.933) 7.939 (3.834) 8.086 (4.450) 8.024 (4.027) 5.517 (3.971) 6.636 (3.680)
Mountain 2 0.366 (0.087) 0.177 (0.078) 0.288 (0.118) 0.409 (0.101) 0.176 (0.095) 0.500 (0.217)
Temple 1 0.511 (0.297) 0.511 (0.302) 0.657 (0.359) 0.789 (0.340) 0.452 (0.284) 0.853 (0.340)
Tiger 0.463 (0.333) 0.571 (0.411) 0.595 (0.430) 0.636 (0.391) 0.413 (0.344) 0.573 (0.392)
Wall 1.708 (0.979) 2.011 (1.231) 1.727 (0.793) 1.692 (0.734) 1.639 (3.210) 2.105 (0.780)
Avg 2.393 (1.015) 1.975 (1.046) 1.943 (1.073) 2.026 (1.016) 2.475 (1.754) 0.995 (2.080)
Avg (w/o Ambush 1) 1.639 (0.952) 1.710 (0.966) 1.750 (1.010) 1.810 (0.951) 1.360 (1.242) 1.727 (0.923)
Table 6: We report the endpoint-error (EPE) on all sequences of Sintel (Butler et al., 2012), shown in the format “final-epe (clean-epe)”. We highlight the best result on each sequence.
@@
We show the sequence-wise results of several representative methods (Morimitsu et al., 2025; Huang et al., 2022; Luo et al., 2024; Saxena et al., 2024; Zhou et al., 2024) on Sintel in Table 6. The sequence ‘Ambush 1’ appears to be an outlier which severely affects the average EPE on the final split. Given the same ImageNet-pretrained Twins backbone, WAFT outperforms SOTA Flowformer++ when ‘Ambush 1’ is excluded.